OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (anonymous):

45=49-4=(7+2)(7-2)= 9 x 5 77= 81-4= (9+2)(9-2)= 11x 7 112= 121-9= (11+3)(11-3)= 14 x 8

OpenStudy (anonymous):

1. (d-7)(d-7) or (d-7)^2 2. (r-7)(r+7) 3. (7b+5)(7b+5) or (7b+5)^2 4. 10u^3 + 4u^2 - 2u + 10 For number 5, I don't quite understand what its asking for ... Do you mean like: a. 45 = (7-2)(7+2) b. 77 = (9-2)(9+2) c. 112 = (11-3)(11+3)